• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 12:42:49 on March10,2005

    Entry number 141088

    keyword=Comment on Gain Settings


    This morning the gains look nicely match, though due to an interesting bug in the calibration code the dE is at a much high gain than the E.

    This was due to the fact there are two different dE energies possible
    for a single E cut. (see figure)

    dE | *
    | *
    | *
    | *
    | *
    --------- E
    ||
    Cut

    After looking at the data we decided to just increase the threshold
    of the dE as the gains were indeed well matched. We did lower the
    HV of all of the dE's by 50 V.